OpenStudy (anonymous):

Please help , Medal & Fan For Best Answer ! <3 <3 A. the sum is (9*10)/2 B. the sum is (9*10) c. the sum is (10*11)/2 d. the sum is (10*11)

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

1=1=(1*2)/2 1+2=3=(2*3)/2 1+2+3=6=(3*4)/2 1+2+3+4= 10= (4*5)/2 1+2+3+4+5=15=(5*6)/2

OpenStudy (anonymous):

What conjecture can you make about the sum of the first 10 integers

OpenStudy (anonymous):

So every time the number goes up once the two numbers you multiply by change once. So 1+2+3+4+5+6= 21 = (6*7)/2 So you would keep going until you to the tenth one. I'll just skip ahead. 1+2+3+4+5+6+7+8+9+10= 55= (10*11)/2